OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with π/10.

OpenStudy (anonymous):

Co-terminal: \[\theta = \theta \pm 2k \pi \] for any integer k

OpenStudy (anonymous):

adding or subtracting multiples of 2pi returns to the same angle

OpenStudy (anonymous):

can you show me how you do that ?

OpenStudy (anonymous):

\[\frac{ \pi }{ 10 } = \frac{ \pi }{ 10 } + 2\pi = \frac{ \pi }{ 10 } + 4\pi = \frac{ \pi }{ 10 } - 14\pi\]

OpenStudy (anonymous):

choose one. and to add fractions you need the same denominator. \[\frac{ \pi }{ 10 } + 2\pi = \frac{ \pi }{ 10 } + 2\pi \frac{ 10 }{ 10 } = \frac{ \pi }{ 10 } + \frac{ 20 \pi }{ 10 } = \frac{ 21\pi }{ 10 }\]

OpenStudy (anonymous):

20π/10 ; -π/10 ?

OpenStudy (anonymous):

neither unfortunately

OpenStudy (anonymous):

a negative one would be:\[\frac{ \pi }{ 10 } - 2\pi = \frac{ \pi }{ 10 } - 2\pi \frac{ 10 }{ 10 } = \frac{ \pi }{ 10 } - \frac{ 20 \pi }{ 10 } = \frac{ -19\pi }{ 10 }\]

OpenStudy (anonymous):

Ohh i'm confused haha

OpenStudy (anonymous):

is it a multiple choice answer? and which part confuses you?